OpenStudy (anonymous):

I need to turn a ratio into a percentage the ratio is 1 out of 18,000 please help

OpenStudy (solomonzelman):

1 / 18,000 use your calculator, and then times 100

OpenStudy (solomonzelman):

0.0055555555....≈0.01

OpenStudy (anonymous):

thank you very much

OpenStudy (solomonzelman):

Anytime, gela!

hero (hero):

to convert a fraction to a percent, multiply it by 100, then add the percent symbol.

hero (hero):

\[\frac{1}{18,000} \times 100 = \frac{1}{180}\] \[\frac{1}{18,000} = \frac{1}{180} \%\]